OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x^3 – 2x^2 – 16x + 32 .

hero (hero):

Hint: Let f(x) = 0 and factor by grouping to get 0 = x^2(x-2)-16(x-2) 0 = (x-2)(x^2 - 16)

OpenStudy (anonymous):

I've got the answers... I think. I came to -4 , 4, and 2. Does that seem right?

hero (hero):

You don't have to "approximate" You can get the actual number

hero (hero):

Yup so 3 real 0 imaginary

OpenStudy (anonymous):

Thank you.